# What Do Temporal Graph Learning Models Learn?

### v2

@@ Appendix C Results with Alternative Hyperparameters
Experiments have already explored the hyperparameter spaces for these models for a range of empirical datasets [undefaa]. As such, we focus our resources on the synthetic data. Nevertheless, given the number of different datasets, seeds and models, extensive experimentation is costly. For hyperparameter tuning, we restricted the generation and training seeds to three variants rather than five.
Following [undefaa], for JODIE we vary the dropout at intervals of 0.1 from 0.0 to 0.6, resulting in 99 additional models trained. For TCL, we switch to using uniform neighbor sampling and vary dropout at intervals of 0.2 from 0.0 to 0.6 (with fewer variants than JODIE due to the larger overall search space).
For persistence, Figure 18 shows the changes in dropout value have very little effect on the predictions from JODIE. With higher values of dropout, it can be seen in Figure 17(a) that TCL stops distinguishing between the two groups of edges, whilst uniform neighbor sampling completely degrades performance. The same effects are observed for periodicity in Figures 17(b) and 19. The recency results for both models do not change with variations of hyperparameters, as shown in Figures 20 and 16.

### v3

@@ Appendix C Results with Alternative Hyperparameters
Experiments have already explored the hyperparameter spaces for these models for a range of empirical datasets [28]. As such, we focus our resources on the synthetic data. Nevertheless, given the number of different datasets, seeds and models, extensive experimentation is costly. For hyperparameter tuning, we restricted the generation and training seeds to three variants rather than five.
Following [28], for JODIE we vary the dropout at intervals of 0.1 from 0.0 to 0.6, resulting in 99 additional models trained. For TCL, we switch to using uniform neighbor sampling and vary dropout at intervals of 0.2 from 0.0 to 0.6 (with fewer variants than JODIE due to the larger overall search space).
For persistence, Figure 16 shows the changes in dropout value have very little effect on the predictions from JODIE. With higher values of dropout, it can be seen in Figure 17(a) that TCL stops distinguishing between the two groups of edges, whilst uniform neighbor sampling completely degrades performance. The same effects are observed for periodicity in Figures 17(b) and 18. The recency results for JODIE, TCL, and TGN do not change with variations of hyperparameters, as shown in Figures 19, 20 and 21.